What mass of PCl5 will be produced from the given masses of both reactants?


1
Expert's answer
2021-02-25T03:20:01-0500

P4 + 10 Cl2 --> 4 PCl5

let the mass of P4 be 20 grams, and the mass of Cl2 be 10 grams

"n(P4)=\frac{m}{M}=\frac{20}{123.896}=0.16"

"n(Cl2)=\frac{m}{M}=\frac{10}{70.906}=0.14"

chlorine in the bottle, we decide further on the chlorine

"\frac{0.14}{10}=\frac{x}{4}"

x=0.056 amount of substance PCl5

"m=n\times M=0.056\times208.239=11.66" g